3. In the model of the steady-state unemployment rate with a fixed labor force, describe the meaning the rate of job finding, the rate of job separation, and the natural rate of unemployment.
Answer #1

The constant unemployment rate in the long run is said to be steady-state unemployment.

Natural rate of unemployment is that rate which is an average estimate around which the unemployment in the economy is fluctuating. This is determined by comparing rate of people finding jobs and rate of job separation. In other words summation of frictional and the structural unemployment is said to be as natural rate of unemployment.

Natural rate of unemployment in steady state can be calculated - U/L (uneployment rate) =s/s+f (s= job searching, f=job finding)

The number of people who finds job is said to be the rate of job finding and rate of job separation is the number of people who quits a job.

Job separation Rate : F*U = S*E ( E- employed, U- unemployed, F- rate of job finding, S- rate of job separation)

Unemployment and Job finding rates exhibit a negative relation whereas, unemployment and job separation exhibits a positive relation). That is, higher S leads to higher unemployment and higher F gives lower unemployment.